OpenStudy (anonymous):

what is the effect of the difference of squares pattern on the multiplication of radicals

OpenStudy (anonymous):

since \[(a+b)(a-b)=a^2-b^2\] then for example \[(2+\sqrt3)(2-\sqrt3)=4-3\]

OpenStudy (anonymous):

and the radical is gone!

OpenStudy (anonymous):

they're conjugates and you end up without a cross (middle) term. same thing with complex numbers

OpenStudy (anonymous):

they can be also known as conjugates

OpenStudy (mathstudent55):

This is used to simplify radical expressions that have a radical in the denominator.